# Refactoring Codebases through Library Design

### v1

@@ 9 Conclusion
We introduce a new benchmark MiniCode and method Librarian for compressing code sources through reusable abstractions. We highlight the challenges that modern models face when producing modular and maintainable code, then present an effective method for using LLMs to do this task in small-scale programs. By framing refactoring as both a design and compression task, our work opens new directions for building more general and scalable code understanding and generation systems. In particular, the structure of MiniCode lends itself well to a reinforcement learning-based task, which synthesizes collections of repositories to refactor then computes rewards based on compression.
Limitations Limitations of this work include the evaluation of synthetic repositories, the poor performance of code agents on refactoring, and the fact that compression may not be correlated with reuse. While our experiments on downstream programming problems partly address the question of reuse, investigating how to measure and encourage reuse for large-scale multi-repo library creation remains an open problem.
Ethics Statement
As with any coding benchmark, our work has the potential to improve automated coding. However, this has the potential to produce malicious code if left to run unchecked. Care should be taken to monitor inputs and outputs to the system.
@@ Acknowledgements
Appendix A Algorithm
@@
4:$\mathcal{C}\leftarrow\text{Cluster}(P_{initial})$
5:$\mathcal{L}_{final}\leftarrow\emptyset$, $P_{final}\leftarrow\emptyset$
@@
7:     $T_{C}\leftarrow\text{GroupIntoTuples}(c)$triangleright Get tuples for each cluster
@@
12:              ($(\{f_{new,i}\},\{\rho^{\prime}_{i}\}\})\leftarrow\textsc{Sample}\left(f_{retrieved},\tau\right)$
@@
15:         $(f_{best},\{\rho^{\prime}_{best}\})\leftarrow\text{RerankAndSelectBest}(S,\ell\left(\cdot\right))$triangleright Rerank using objective
@@
A.1 Objective Functions
Human evaluation best aligns with the MDL objective function.
We collect 19 CodeContests problems and a pair of refactorings for each, and have 4 human judges pick which they prefer. Each refactoring comprises three rewritten programs, plus the library generated by Librarian. Each pair of refactorings passed all testcases, and were chosen to include both the highest and lowest perplexity refactorings. Human judges are authors on this paper, but were ‘blinded’ to not know relative objective function scores for the candidate refactorings.
An MDL objective aligns best with human preferences (Figure 5), although number of tokens is also a reasonable proxy for human judgment. That token count and MDL yield different statistics furthermore implies that the lowest MDL program is not always the shortest program.
Figure 5: Human evaluation of different refactoring objectives. Humans and models give preferences to pairs of refactorings/libraries.
@@ A.1 Objective Functions
Appendix B Collection Coherence Measures
We provide the full definitions of the collection coherence measures.
@@ Appendix B Collection Coherence Measures
Appendix C Benchmark Comparison
We compare our benchmark, MiniCode, to similar benchmarks in Table 4. We define creativity and design as the need to explore diverse solutions in order to find the best solution possible. For example, optimizing for program correctness alone does not require exploring a large solutions space, whereas optimizing a program for speed would. In the case of compressing large code sources, we must explore the large space of shared abstractions afforded by libraries in order to maximize compression.
@@ Appendix C Benchmark Comparison
SWE-bench [9] Low Repository
Commit-0 [10] Medium Repository
RefactorBench [11] Low File
ECCO [13] High Function
KernelBench [14] High Function
MiniCode(Ours) High Multi-repository
Appendix D Full MiniCode Repository results
We present the full repository-level results of MiniCode-repository in Tables 5 and 6.
Collection Agent LLoC CC MDL ratio Pass rate
datapipe original 474 116 1.0 1.0
Cl-Cl 1084 341 9.1 1.0
Cl-Cx 645 174 2.2 1.0
state_machine original 619 222 1.0 1.0
Cl-Cl 2271 735 6.3 0.9
Cl-Cx 686 227 1.9 1.0
config_schema original 949 284 1.0 1.0
Cl-Cl 922 339 2.5 failed
Cl-Cx 626 207 2.1 1.0
cli_tools original 875 300 1.0 1.0
Cl-Cl 2925 909 4.0 1.0
Cl-Cx 5848 2378 3.1 failed
cli_form original 352 174 1.0 1.0
Cl-Cl 855 342 3.8 1.0
Cl-Cx 1366 548 2.9 1.0
Table 5: Full results on MiniCode-repositories small.
Collection LLoC CC MDL ratio Pass rate
file_system_analyzer original 3911 1338 1.0 1.0
refactored 6495 2218 1.6 1.0
query_language_interpreter original 5181 2105 1.0 1.0
refactored 7193 2966 1.4 1.0
personal_knowledge_management original 5341 1832 1.0 1.0
refactored 6357 2364 1.4 1.0
virtual_machine_emulator original 6841 2283 1.0 1.0
refactored 10108 3580 1.6 1.0
personal_finance_tracker original 4306 1482 1.0 1.0
refactored 7443 2512 1.8 0.6
in_memory_database original 4565 1671 1.0 1.0
refactored 6667 2556 1.5 1.0
command_line_task_manager original 7964 2921 1.0 1.0
refactored 10387 3899 1.3 1.0
text_editor original 4324 1323 1.0 1.0
refactored 5792 1822 1.2 1.0
concurrent_task_scheduler original 10350 3935 1.0 1.0
refactored 18653 7165 1.9 0.8
incremental_backup_system original 4587 1482 1.0 1.0
refactored 5365 1788 1.3 failed
Table 6: Full results on MiniCode-repositories large.
Appendix E Data Generating Prompts
Prompt to generate general library text descriptions.
1I need ideas for Python libraries that can be implemented by language models. These libraries should:
2
31. Be implementable using only Pythons standard library - no external dependencies
42. Have enough complexity to demonstrate sophisticated code design (10-20 functions/methods)
53. Include room for interpretation, so that different implementations can be unique while sharing core functionality
64. Have clear, practical utility that solves a real programming need
75. Be realistically implementable by an intelligent language model
86. Be testable with pytest
97. Include opportunities for different design approaches (functional vs OOP, etc.)
10
11For each library, provide a description that outlines:
12- The problem domain and core purpose
13- Key required functionality (without being too prescriptive about implementation details)
14- Potential use cases that demonstrate practical applications
15- Suggested extension points where implementers could add their creative spin
16
17Please generate several proposals in markdown, following this format:
18
19‘‘‘file:<library_name>/DESCRIPTION.md
20# <Library Name>
21
22## Purpose and Motivation
23<3-5 sentences on what problem this library solves and why its useful>
24
25## Core Functionality
26<Description of 4-6 high-level key features/capabilities without specifying exact implementation>
27‘‘‘
28
29Be creative! Focus on domains where standard Python libraries provide enough building blocks but where a well-designed abstraction layer would add significant value.
Listing 2: Prompt to generate library descriptions
1
2Consider a code repository designed to support the following task description:
3‘‘‘
4{task_content}
5‘‘‘
6
7Please list a couple dozen features that would be useful for the repository. Include the specified features as well as several others.
8
9List the suggested feature names and descriptions in the following format:
10
111: <feature_1_name>: <feature_1_description>
122: <feature_2_name>: <feature_2_description>
133: <feature_3_name>: <feature_3_description>
14...
1530: <feature_30_name>: <feature_30_description>
16"""
17
18persona_prompt_template = """Consider the following features of a code repository:
19{listed_features}
20
21Think about several possibilities for what kind of person might use this code repository and what they might use it for. Please write several brief descriptions for the code repository in first person, formatted in markdown as follows:
22‘‘‘file:{library_name}/<persona_name>/TASK.md
23# The Task
24
25I am a <...> I want to be able to <...> This code repository <...>
26
27# The Requirements
28
29* ‘<function_name>‘ : <feature description>
30* ...
31‘‘‘
32
33Be creative! Write the task description in the style of the proposed persona. Be as exhaustive as possible in including the listed features in the task descriptions requirements.
Listing 3: Prompt to generate potential uses and features given a library description
1
2
3I need you to implement a Python solution and COMPREHENSIVE suite of tests based on the following task files.
4Your code must pass the tests provided.
5
6{task_content}
7
8CRITICAL FORMATTING INSTRUCTIONS:
91. You MUST format ALL code files exactly as shown below - no exceptions
102. Start each file with the markdown codeblock marker, followed by "file:" and the relative path
113. End each file with the closing markdown codeblock marker
124. Do not use any other format or markdown variations
135. For test files, do not put them in a subdirectory-- keep them in the outermost level.
14
15For each source code file:
16
17‘‘‘file:<relative_file_path>
18<file_content>
19‘‘‘
20
21For each test file:
22
23‘‘‘file:<relative_file_path starting with test_>
24<test_file_content>
25‘‘‘
26
27IMPORTANT:
28- The opening format must be exactly: ‘‘‘file:path/to/file.py
29- Do not add language indicators like ‘‘‘python
30- Do not add explanations between files
31- Each file must be contained within its own codeblock with the precise format shown above
32- The system parsing your response requires this exact format to function properly
33
34EXAMPLE OUTPUT FORMAT:
35‘‘‘file:mymodule/mymodule.py
36def example_function():
37 return "This is a sample function"
38‘‘‘
39
40‘‘‘file:test_utils.py
41import pytest
42from mymodule.mymodule import example_function
43
44def test_example_function():
45 assert example_function() == "This is a sample function"
46‘‘‘
47
48Begin your implementation now, following these formatting rules precisely.
Listing 4: Prompt to generate initial implementation
1
2I need you to fix the implementation of the following code that is failing tests.
3
4# Current Implementation:
5{src_code_content}
6
7# Test Files:
8{test_content}
9
10# Failed Tests:
11{failed_test_details}
12
13# Test Output (if available):
14{test_output}
15
16Please carefully analyze the errors and test failures. Pay special attention to:
171. The exact assertion failures or error messages
182. What the tests expect vs. what your current implementation provides
193. Any edge cases or special conditions you might have missed
20
21Your task is to fix the implementation to make all tests pass. For each file that needs to be modified,
22provide the content in the following format:
23
24‘‘‘file:<relative_file_path>
25<file_content>
26‘‘‘
27
28Where <relative_file_path> is the relative path to the file and <file_content> is the updated content of the file.
29Focus on fixing the specific issues identified in the errors and failed tests while maintaining the overall structure of the code.
30
31IMPORTANT: Make targeted changes to address the specific failing test cases. Make sure your implementation passes all test cases,
32including any edge cases or special conditions mentioned in the tests. Be sure to output code in the specified format.
Listing 5: Prompt to fix code implementation, given pytest output.
Appendix F Refactoring examples of Librarian on Code Contests
F.1 Example 1
@@
Listing 6: Version 1, New Helpers
@@ 30 O = [trie_pop_min_xor(trie, a) for a in A]
Listing 7: Version 1, Refactored Programs
@@
Listing 8: Version 2, New Helpers
@@
Listing 9: Version 2, Refactored Programs
F.2 Example 2

### v3

@@ 8 Conclusion
We introduce a new benchmark MiniCode and method Librarian for compressing files through reusable abstractions. We highlight the challenges of producing modular and maintainable libraries, then present an effective method for using LLMs to do this task. By framing refactoring as an optimization problem, our work opens new directions for building more general and scalable code understanding and generation systems. In particular, the structure of MiniCode lends itself well to reinforcement learning, where training would entail synthesizing collections of repositories to refactor then computing rewards based on MDL or other metrics.
Limitations. We evaluate on synthetic toy problems (Logo and Date), and on competition programming problems, neither of which are naturalistic, although this is partly counterbalanced by our study of real-world refactoring. Compression, whether measured by tokens or MDL, may not always correlate with reuse, a limitation we sought to address through our experiments on downstream programming problems, and on holdout Transformers files: But investigating reuse on unseen app-building problems for real-world repo-level refactors remains open.
@@ Acknowledgements
KE and ZC were supported by NSF grant #2310350.
Appendix A Incremental Version of Librarian
We support sharing library components across clusters through the below incremental version of our algorithm. It processes clusters sequentially, rather than in parallel, and its output can depend on the ordering imposed upon the clusters.
$\mathcal{L}_{0}=\varnothing$
$\{\rho_{n}^{\prime}\}=\bigcup_{c\in\text{CLUSTER}(\{\rho_{n}\})}\{\rho_{c,i}^{\prime}\}_{i=1}^{|c|}$
$(\Delta\mathcal{L}_{c},\{\rho_{c,i}^{\prime}\}_{i=1}^{|c|})=\arg\min_{(\Delta\mathcal{L},\{\rho_{i}^{\prime}\})\in\text{SAMPLE}_{k}(c;\,\mathcal{L}_{t-1})}\ \ell(\mathcal{L}_{t-1}\cup\Delta\mathcal{L},\{\rho_{i}^{\prime}\}_{i=1}^{|c|})\quad\forall c\in\text{CLUSTER}(\{\rho_{n}\})$
$\mathcal{L}_{t}=\mathcal{L}_{t-1}\cup\Delta\mathcal{L}_{t},\qquad\mathcal{L}^{\star}=\mathcal{L}_{|\text{CLUSTER}(\{\rho_{n}\})|}$
Appendix B Algorithm
@@
[…]
@@
7:  $T_{C}\leftarrow\text{GroupIntoTuples}(c)$ $\triangleright$ Get tuples for each cluster
@@
12:     $(\{f_{new,i}\},\{\rho^{\prime}_{i}\}\})\leftarrow\textsc{Sample}\left(f_{retrieved},\tau\right)$
@@
15:   $(f_{best},\{\rho^{\prime}_{best}\})\leftarrow\text{RerankAndSelectBest}(S,\ell\left(\cdot\right))$ $\triangleright$ Rerank using objective
@@
Appendix C Experimental Setup
Grouping Programs into Collections
To facilitate parallel application of Librarian and manage the dataset scale, we assume that semantically distant files will have minimal overlap in their optimal library functions. Therefore, our overall approach partitions the dataset into disjoint collections through clustering.
For CodeContests, these collections are constructed from an initial corpus of $\sim$9k problems with Python solutions: We first filter these files, removing those whose selected canonical solution is under 10 lines (minimal refactoring potential). For the remaining 4596 solutions we use a language model to generate textual descriptions of canonical solutions—emphasizing reusable components—which are embedded using OpenAI’s text-embedding-ada-002.
Agglomerative Clustering (Ward Jr, 1963) is subsequently applied to these embeddings to partition the files into a predefined number of initial clusters, in our case 120. To create uniformly sized experimental units, we subsample each such cluster to form collections of 30 files. This collection size was empirically chosen because it balanced between the runtime of Librarian without limiting compression. We select 10 collections that we then use to evaluate our methods.
For Transformers, since the number of models is on the lower end, we manually chose a set of popular LLM / VLM models and passed them to the agent in collections of 5 code sources.
REGAL Baselines.
To evaluate the ability of our libraries to support reuse on new problems, we turn to the program synthesis tasks used in REGAL, where learned libraries are added to help the program synthesizer. We evaluate on the two domains published by the authors, Logo and Date. Because our clustering is inspired by REGAL but adds additional complexity, for fair comparison, we keep their setup the same and only augment the training using sample + MDL rerank procedure described in Section˜5.
Code Contests.
To evaluate Librarian on refactoring Code Contests we select 6 collections of 30 files (problems). In each collection we group the problems into tuples of size $3$. We set the sample budget to be $K=8$, since our ablations show that with larger $K$ we discover better libraries 2. We use the MDL objective for rankings. The model used for sampling is OpenAI’s o4-mini (OpenAI, 2024). To obtain MDL scores we use Qwen 2.5 7B Instruct (Qwen et al., 2025) as a balance between quality, speed, and cost.
Code Agents on Transformers and Diffusers Repositories.
To fairly evaluate performance on the task by state-of-the-art systems, we use coding agents that advertise long-context ability to reason about, write, and refactor code repositories. Specifically, we use Claude Code (Cl) (Anthropic, 2025) which uses the Opus 4.1.
We test whether code agents can refactor collections of code sources autonomously, without human intervention. Refactoring repositories with code agents involves planning and iterative (re-)implementation and testing. Code agents are prompted to perform each of these steps, with feedback from the unit tests. Agents must run and repair unit tests autonomously. We run coding agents multiple times per task, logging their progress in checklists stored in text files.
@@ Appendix C Experimental Setup
Appendix D Further Results on Refactoring Metrics
Here we include the full graph of asymptotic behavior of all four scoring metrics (MDL, tokens, MI and CC), reporting the raw library metrics and the metric ratios. We can see that refactored programs have higher CC than the baseline and that optimizing CC as an objective does not decrese the other metrics. MDL performs best on total raw library usage as well as on average times a library function is used in the refactorings. MI ends up optimizing the number of library functions the best, but their reusability is below the average for the refactorings. Optimizing for tokens produces smaller libraries with less usage per function compared to optimizing MDL.
Figure 7: Asymptotic behavior of metrics for scoring libraries and refactorings (columns) varying refactoring budget (horizontal axes).
Appendix E Human Study Details
We ran a user study with 14 participants where each participant had to judge 10 refactoring pairs. Each pair was comparing two out of three metrics, MDL, tokens, and MI. We showed the participants the original programs, as well as both of the refactoring versions (including the programs and the learned libraries). The participants were able to choose either version or say that the two refactorings are almost the same.
To quantify pairwise preferences between refactoring metrics, we employed the Bradley-Terry model, a standard framework for analyzing paired comparison data. We fit the model using maximum likelihood estimation with mutual information (MI) as the reference category $(\pi_{M}I=1)$. To address potential noise in human judgments, we applied consensus-based filtering with a 75% threshold, retaining all responses for comparisons with low consensus (indicating genuine ambiguity) while excluding minority responses on high-consensus comparisons where the majority preference likely indicates the correct judgment. This conservative approach preserved 92.4% of responses while strengthening the statistical evidence for metric preferences, with MDL significantly outperforming MI ($p=0.67$, $95\%$ Confidence Interval: $[0.56,0.78]$). Even without the filtering MDL preference over MI was statistically significant.
Appendix F Best@k Compression is a U-Statistic
We wish to estimate the expected compression ratio achieved by our sample + rerank method, which samples $k$ candidate refactorings, discards any that do not pass the tests, and selects the one with the lowest score (total log-prob).
Background on U-Statistics. Let $Z_{1},\dots,Z_{n}\overset{\text{i.i.d.}}{\sim}F$. For a symmetric function $h:\mathcal{Z}^{k}\to\mathbb{R}$, the U-statistic of order $k$ is defined as
$U_{n}\;=\;\binom{n}{k}^{-1}\sum_{1\leq i_{1}<\cdots<i_{k}\leq n}h(Z_{i_{1}},\dots,Z_{i_{k}}).$ (6)
By construction,
$\mathbb{E}[U_{n}]\;=\;\mathbb{E}[h(Z_{1},\dots,Z_{k})],$ (7)
so $U_{n}$ is an unbiased estimator of the population quantity $\theta=\mathbb{E}[h(Z_{1},\dots,Z_{k})]$.
Application to Best@k Compression. Let each valid refactoring be a pair $Z=(S,C)$, where $S$ is the score and $C$ is the compression ratio. Define the symmetric function
$h_{k}(z_{1},\dots,z_{k})\;=\;C_{j^{*}},\qquad j^{*}=\arg\min_{1\leq j\leq k}S_{j},$ (8)
the compression ratio of the lowest-score refactoring among $k$ draws. The population target is then
$\theta_{k}=\mathbb{E}[h_{k}(Z_{1},\dots,Z_{k})].$ (9)
Given $n$ valid samples, our estimator is
$\widehat{\theta}_{k}\;=\;\binom{n}{k}^{-1}\sum_{1\leq i_{1}<\cdots<i_{k}\leq n}h_{k}(Z_{i_{1}},\dots,Z_{i_{k}}).$ (10)
Proposition. $\widehat{\theta}_{k}$ is a U-statistic of order $k$ with function $h_{k}$, and hence an unbiased estimator of $\theta_{k}$.
Proof.
(1) Symmetry of the function $h_{k}$. $h_{k}$ selects the compression associated with the lowest score among its $k$ arguments. Permuting the inputs does not affect this outcome (ties can be resolved with a fixed, permutation-invariant rule). Thus $h_{k}$ is symmetric.
(2) U-statistic form. By definition, a U-statistic of order $k$ with kernel $h_{k}$ is
$U_{n}=\binom{n}{k}^{-1}\sum_{1\leq i_{1}<\cdots<i_{k}\leq n}h_{k}(Z_{i_{1}},\dots,Z_{i_{k}}),$
which matches $\widehat{\theta}_{k}$ exactly.
Therefore, $\widehat{\theta}_{k}$ is a U-statistic of order $k$. By the unbiasedness property of U-statistics,
$\mathbb{E}[\widehat{\theta}_{k}]=\theta_{k}.$
$\square$
Thus, our reported best@k compression curves provide unbiased estimates of the expected performance of the sample + rerank method.
Appendix G Clustering Analysis: CodeContests
We analyze the coherence of the clusters underlying collections in MiniCode-CodeContests. In particular, we compare clustering based on o4-mini generated file descriptions against task descriptions. Since task descriptions in competition coding problems are designed to hide the algorithmic approach needed to solve problem, we expect that clusters based on file descriptions are more coherent. We use Normalized Tag Instance Entropy and Herfindahl-Hirschman Index to evaluate clusterings. Figure˜8 shows our clustering approach yields more thematically coherent clusters, evidenced by achieving lower entropy and higher HHI values across the entire tested range of $N$. We provide definitions of our measures below.
G.1 Collection Coherence Measures
We use two measures to evaluate the thematic coherence of collections: Good collections should group files with a (1) concentrated and (2) identifiable set of shared conceptual tags, which for CodeContests are provided as ground truth (trees, graphs, etc.).
We provide the full definitions of the collection coherence measures below.
@@ G.1 Collection Coherence Measures
Appendix H Benchmark Comparison
We compare our benchmark, MiniCode, to similar benchmarks in Table 4. We define creativity and design as the need to explore diverse solutions in order to find the best solution possible. For example, optimizing for program correctness alone does not require exploring a large solutions space, whereas optimizing a program for speed would. In the case of compressing large files, we must explore the large space of shared abstractions afforded by libraries in order to maximize compression.
@@ Appendix H Benchmark Comparison
SWE-bench (Jimenez et al., 2024) Low Repository
Commit-0 (Zhao et al., 2025) Medium Repository
RefactorBench (Gautam et al., 2025) Low File
ECCO (Waghjale et al., 2024) High Function
KernelBench (Ouyang et al., 2025) High Function
MiniCode(Ours) High Repository
Appendix I Full MiniCode Codecontests Results
We present the full agent scores for the CodeContests split in Table 5. The results are given both for each cluster of code sources, as well as averaged across clusters.
Cluster Agent Tokens $C$ Pass % MDL MDL %
0 original 9088 95 80.3 11745.85 100.0
sonnet 3.7 18114 176 87.0 15005.18 127.7
sonnet 4 11121 138 80.3 9901.53 84.3
codex-mini 9321 95 80.3 9990.74 85.1
1 original 12531 255 89.7 13431.86 100.0
sonnet 3.7 10470 239 96.7 8933.65 66.5
sonnet 4 11325 298 96.7 8214.42 61.2
codex-mini 12762 255 89.7 11798.73 87.8
2 original 14087 376 89.0 15012.77 100.0
sonnet 3.7 17345 429 91.3 13145.02 87.6
sonnet 4 14270 356 93.0 10522.66 70.1
codex-mini 14318 376 89.0 13273.81 88.4
3 original 14261 246 90.3 13348.82 100.0
sonnet 3.7 20749 241 97.7 15859.02 118.8
sonnet 4 13433 197 80.7 11937.04 89.4
codex-mini 14495 246 90.3 11616.41 87.0
4 original 17693 336 80.7 14665.16 100.0
sonnet 3.7 29860 358 100.0 20666.52 141.0
sonnet 4 18684 352 82.0 12801.21 87.3
codex-mini 17923 336 80.7 12902.09 88.0
5 original 12588 286 92.0 12790.11 100.0
sonnet 3.7 10580 128 99.3 8435.12 65.9
sonnet 4 10416 155 99.3 9167.85 71.7
codex-mini 12819 286 92.0 11086.19 86.7
6 original 11020 131 54.3 13540.41 100.0
sonnet 3.7 21747 502 88.0 19446.07 143.6
sonnet 4 11177 143 57.3 10492.00 77.5
codex-mini 11251 131 54.3 11651.65 86.1
7 original 12301 180 80.0 12393.73 100.0
sonnet 3.7 16390 166 91.0 13371.59 107.9
sonnet 4 11625 150 85.7 9304.25 75.1
codex-mini 12534 180 80.0 10549.04 85.1
Avg original 12946 238 82.0 13366.09 100.0
sonnet 3.7 18157 280 93.9 14357.77 107.4
sonnet 4 12756 224 84.4 10292.62 77.1
codex-mini 13178 238 82.0 11608.58 86.8
Table 5: Comparison of the pass rate and compression metrics of the original files, Claude Sonnet 4 and codex-mini refactorings across CodeContests clusters.
Appendix J Refactoring examples of Librarian on Code Contests
J.1 Example 1
@@ Appendix I Full MiniCode Codecontests Results
Listing 2: Version 1, New Helpers
@@ 30 O = [trie_pop_min_xor(trie, a) for a in A]
Listing 3: Version 1, Refactored Programs
@@
Listing 4: Version 2, New Helpers
@@
Listing 5: Version 2, Refactored Programs
J.2 Example 2